OpenStudy (anonymous):

Someone help me please oh my god I'm freaking out. A landscaper wants to plan a walkway that passes between a tree and the border of the lawn. Using these as the focus and directrix, how can the landscaper plan a parabolic path that will be equidistant from the tree and the border at all times? Describe your method in full sentences.

hero (hero):

At @ashaboo456, do you have any question in particular about this problem you have posted?

OpenStudy (anonymous):

Im confused about pretty much everything, ive been on this question for almost an hour now ugh. Sorry, do you mind helping?

hero (hero):

Do you remember that distance formula I gave you for finding the equation of a parabola given the focus and directrix?

OpenStudy (anonymous):

Uh, no, sorry omg

hero (hero):

If given the focus \((x_1, y_2)\) and directrix \((x_2, y_2)\), you can insert both points into this distance formula: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\) Afterwards, you simplify the equation until you have the form: \(y = a(x - h)^2 + k\) or the form \(x = a(y - h)^2 + k\)

OpenStudy (anonymous):

So is that it then? Sorry I feel really stupid right now

OpenStudy (anonymous):

because it doesn't give me any points

hero (hero):

What I gave you is a general formula. To do this problem, you'll have to make up your own points.

OpenStudy (anonymous):

Oh, alright. I think I've got it now. Thank you so much seriously you are a grade saver !